OpenStudy (piercetheveil47):

MEDAL AND FAN. Peter and Jacob are members of different video game libraries. Peter pays a membership fee of $45, and he pays $6 for every video game he rents. The following function shows the total amount of money, y, in dollars, that Jacob pays for renting x number of video games. y = 5x + 25 How many more dollars does Peter pay for a membership fee than Jacob? $11 $20 $25 $45

OpenStudy (piercetheveil47):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the membership fee that Jacob pays

OpenStudy (piercetheveil47):

45?

jimthompson5910 (jim_thompson5910):

that's for peter

OpenStudy (piercetheveil47):

he pays 25, right?

OpenStudy (piercetheveil47):

so the answer is C, 20?

OpenStudy (piercetheveil47):

*B

jimthompson5910 (jim_thompson5910):

yeah as a membership fee and yes, the difference is $20

OpenStudy (piercetheveil47):

can you help me with another one that i really don't understand??

jimthompson5910 (jim_thompson5910):

so if the games are identical (for peter and jacob), then peter is getting ripped off because he pays higher membership fees and pays more money per game to rent

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (piercetheveil47):

Which equation represents a linear function? Equation 1: y = 2x + 1 Equation 2: y2 = 3x + 1 Equation 3: y = 5x5 − 1 Equation 4: y = 4x4 − 1 Equation 1 Equation 2 Equation 3 Equation 4

OpenStudy (piercetheveil47):

those little question marks after 5 and 4 in equation 3 and 4 are negatives

jimthompson5910 (jim_thompson5910):

linear equations are equations that can be written into the form y = mx+b m and b are fixed numbers

OpenStudy (piercetheveil47):

A?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (piercetheveil47):

oh! okay :)